To test the performance of its tires, a car
travels along a perfectly flat (no banking) circular track of radius 368 m. The car increases
its speed at uniform rate of
at ≡d |v|dt = 3.3 m/s2
until the tires start to skid.
If the tires start to skid when the car reaches
a speed of 31.5 m/s, what is the coefficient of
static friction between the tires and the road?
The acceleration of gravity is 9.8 m/s2
1
Expert's answer
2019-12-23T11:41:19-0500

The Newton's second law states


"ma=F_{\rm net}"

In this case


"ma=F_{\rm fric}=\mu N=\mu mg"

The total acceleration


"a=\sqrt{a_n^2+a_{\tau}^2}=\sqrt{\left(\frac{v^2}{R}\right)^2+a_{\tau}^2}"

Hence


"a=\mu g"

Finally, the coefficient of static friction


"\mu=\frac{\sqrt{\left(\frac{v^2}{R}\right)^2+a_{\tau}^2}}{g}"

"=\frac{\sqrt{\left(\frac{31.5^2}{368}\right)^2+3.3^2}}{9.8}=0.435"